Summary

Due to the remodeling of the north half of the 5th floor by the Graphics Group, the 530 Soda machine room is being demolished. Space for existing equipment is being prepared in other parts of Soda, primarily 420A Soda.

Support for this project is being provided by:

  • EECS Facilities
  • Par Lab, RAD Lab, and Lo Cal Technical Staff
  • EECS Networking

Scheduling will be displayed here as will list of outstanding tasks with taskholders.

Registering CS/EECS Systems in 420A

There is now a rack for hosts in CS/EECS.Berkeley.EDU moving from 530 to 420A with EECS Networking provided by the RAD Lab who has graciously provided one of their switches to EECS for this purpose until longer term plans are made. Due to the manner in which the switch is connected to the EECS Network, all system registrations using that switch should use the CableID below

  • Rack: 20
  • CableID: 4B-085-C

Goals

  • Minimal downtime both in duration and frequency:
    • max of 1 day for any system
    • no more than 1 instance for any system - up-down-up, not up-down-up-down-up-down-…-up
      • Systems may get rebooted often during migration, but once we say something is “up”, it should stay up modulo end-user desires/needs.
    • Minimize impact on project/publication deadlines (sooner is most often better)
      • after May 12 for PSI cluster (target May 16?)
      • May 23: tape-out for ParLab/RAMP
      • May 23-25: AMP Lab retreat
      • Jun 1-3: ParLab/LoCal Retreats
      • June 2: for PSI Cluster users
  • Progress towards more well managed machine room space in 420A as a model for other such spaces.

Schedule

  • 2 May - meeting with RAD Lab / Par Lab / LoCal staff (“420A Staff”), EECS Facilities
  • 9 May - followup meeting to above
  • 22 to 31 May - RAD Lab / Par Lab / LoCal staff away at retreats
  • 25 May - power shutdown per Scott
  • 2 Jun 2011 - 530 Soda must be ready for contractor demolition for Graphics Group Remodel

Tasks

  • Confirmed: that we can use the DETER rack from 190 Soda - used by PSI
  • 9 May, DaveS, power survey of 420A to determine location, type, and status of underfloor power-taps.
    • Eventually, DaveS, get power survey to everyone, esp JonK to put up online.
  • 9 May, 420A Staff, tagging of all equipment in 420A to “Keep”, “E&S”, or “Move to (Location)”
  • ASAP, JonK, coordinate with EECS networking to prep for network changes in 420A.
    • Jeff has initiated conversation with <networks@eecs>: [Ticket #17733] support@millennium
    • <networks@eecs> working on network changes - should be ready on/by 9 May –JonK
      • Update, <networks@eecs> won't have switch ready until 10 May –JonK
      • Update, <del>JonK to get switch config(s) to Janet/<networks@eecs> Done 11 May</del>
      • Update, Janet/<networks@eecs> punting to Albert, eta unknown
    • Jon to configure networking equipment - 17-19 May Done
  • 9 May, 420A Staff, list power needs for incoming equipment (circuit size, connector, location) - ongoing
  • 9 May, JonK needs to prep a new rack for rackslot 23
    • Would like a 2nd L6-30R there, but can deal with the L6-30R/L5-30R combo.
  • 10 May, JonK, Clear RelayRack 36 for Demo for new PSI Rack
  • 16 May, JeffA/JonK, target move date for PSI cluster nodes

Move Schedule

  • Anytime
    • batoka (moved-down)
    • maramba (moved-down)
    • ontario (moved-down)
    • quickia (UP)
    • galaxy (moved-UP)
    • gainestown (moved-UP)
    • emerald (moved-UP)
    • cuda1/2 (moved-UP/moved-UP)
    • www-research2.cs (moved-UP)
    • huron
    • r58 (moved-UP)
    • doradi (moved-UP)
    • dhruva.cs (moved-UP)
    • rx1/2
  • 9 May - s141-s144 *done*
  • > 10 May - mustang(moved/UP)/taurus(moved/UP)
  • > 16 May - PSI (s85-s104:moved-UP)
    • s105-108 (s105-s107:up; s108:down)
  • Unscheduled
    • RAD/Par/PSI
      • parcad
      • beckton
      • boxboro
      • knight
      • ferry
      • r1-r5
      • bee need to finish rewiring to front-end host
      • thumper
      • scc/boathouse/boathouse2
      • kubi2.cs
      • spl.cs
      • some-psi box moved to storage
      • texas to be brought up, time TBD
      • x51/x52 moved to storage
    • Graphics
      • brunswik
      • glamdring
      • menegroth
      • balin
      • bifur
      • bofur
      • bombur
      • thorin
      • gloin
      • qssc-s4r moved by CUSG to 165 Cory to moved back, schedule TBD
      • some-dell joubert
    • BMRC - <fraser@eecs>
      • cls1 (rowe/bmrc) Handled by <fraser@eecs>
      • quimby to be wiped and donated to Theory Group
      • gumby to be wiped and donated to Theory Group
    • <dawnsong@cs>
      • aerie.cs pending System Update by <kevinchn@eecs>
      • arkbrid.cs pending System Update by <kevinchn@eecs>
      • bullseye.cs pending System Update by <kevinchn@eecs>
      • orchid.cs pending System Update by <kevinchn@eecs>

Network Notes

There are a total of 21 *.cs hosts potentially being moved from 530 to 420A.

As mentioned above, CS/EECS hosts may obtain a network port in 420A via the switch located in the top of rack 20, CableID: 4B-085-C.

RAD / Par Lab systems

  • Current plan for Par/RAD Lab systems is to daisy-chain three more Cisco 4948-10GE switches into existing DCOD (Daisy Chain Of Doom): 1 PSI from 530, 1 existing Par/RAD from 530, 1 new Par/RAD
  • Eventual plan is to have a loop terminating on both ends at the 288 Cisco 6506-E with STP handling loop detection and link failure detection/mitigation.
  • kubi2.cs [currently on 5B-082-C]
  • spl.cs [currently on 5B-086-C]
  • www-research2.cs [currently on 5B-086-B]
  • dhruva.cs

Graphics Systems

  • 14 *.cs machines belonging to job, malik, roc

Dawn Song's Systems

  • aerie.cs
  • arkbird.cs
  • bullseye.cs
  • orchid.cs
 
530-migration.txt · Last modified: 2011/05/24 10:31 by jonah
 
Except where otherwise noted, content on this wiki is licensed under the following license:CC Attribution-Noncommercial-Share Alike 3.0 Unported
Recent changes RSS feed Donate Powered by PHP Valid XHTML 1.0 Valid CSS Driven by DokuWiki